if 27 ia added to a two digit number the result is a number with the same digit, but in reverse order. the sum of the digits is 11
Answer by mananth(16946)

Let the number be xy
x in the tens place and y in units place

1 x + 1 y = 13 .............1
-9 x + 9 y = 27 .............2
9 y
multiply (1)by 9
Multiply (2) by -1
9 x 9 y = 117
9 x + -9 y = -27
Add the two equations
18 x = 90.000
/ 18
x = 5
plug value of x in (1)
1 x + 1 y = 13
5 + 1 y = 13
1 y = 13-5
1 y = 8
y = 8